A rectangular room is 9 feet longer than twice the width. If the room's perimeter is 186 feet, what are the room's dimensions?
Answer by Earlsdon(6294)

Starting with the perimeter: P = 2(L+W) and this is 186 feet, so:
186 = 2(L+W) Dividing both sides by 2:
93 = L+W But:
L = 2W+9 Substitute into the first equation.
93 = (2W+9)+W
93 = 3W+9
3W = 84
W = 28 and:
L = 2(28)+9
L = 56+9
L = 65
Check:
2(L+W) = 2(65+28) = 2(93) = 186.